OpenStudy (anonymous):

2. Find the slope of a line that passes through (-2,-3), and (1, 1) . a. 1/3 b. 2 c. 2 d. 4/3

OpenStudy (anonymous):

@mathmath333 please help

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

O_O

OpenStudy (texaschic101):

are you lost or do you understand hartnn's explanation ??

OpenStudy (anonymous):

I'm lost still. (I'm an idiot )

OpenStudy (anonymous):

:(

OpenStudy (mathmath333):

\(\large\tt \color{black}{(x_1,y_1)=(-2,-3)}\) \(\large\tt \color{black}{(x_2,y_2)=(1,1)}\)

OpenStudy (texaschic101):

you will be using the slope formula : slope = (y2 - y1) / (x2 - x1) now you have 2 sets of points set 1 (x1, y1) and set 2 (x2,y2) now label your points... (-2,-3)....x1 = -2 and y1 = -3 (1,1)...x2 = 1 and y2 = 1 now we sub slope = (1 - (-3) / (1 - (-2).....now solve...but remember that a negative timed=s a negative is a positive. Can you do it now ?

OpenStudy (texaschic101):

typo....a negative times a negative is a positive